Database last updated on Mon, 02 Feb 2026 00:40:50 +0000 by localhost

Instrument constant changes for the week before 2023309 (1 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 16 rows)

Measurements and tests summary.


_jday_loc_serial_skcdtedfmhgrssluv
309RESOLUTE#031ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.198)1247201643
(0 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B30923.0311#0313534